OpenStudy (anonymous):

HEEEELLLLP- solve for x x-4\left[\begin{matrix}2 & -8 \\ -4 & 2\end{matrix}\right]=\left[\begin{matrix}4 & -6 \\ 2 & -8

OpenStudy (anonymous):

x-4\left[\begin{matrix}2 & -8 \\ -4 & 2\end{matrix}\right]=\left[\begin{matrix}4 & -6 \\ 2 & -8

OpenStudy (anonymous):

x-4\left[\begin{matrix}2 & -8 \\ -4 & 2\end{matrix}\right] =\left[\begin{matrix}4 & -6 \\ 2 & -8